OpenStudy (anonymous):

I am...confused. The average (arithmetic mean) of n and 3 is equal to the average of 7 and 5. What is the value of n ? (A) 7 (B) 8 (C) 9 (D) 10 (E) 12

OpenStudy (amistre64):

equate the two and solve for n

OpenStudy (mathstudent55):

How do you find the average of two numbers?

OpenStudy (anonymous):

you add them...then divide by the amount of numbers added. (I think)

OpenStudy (mathstudent55):

Right. Add the numbers and divide by 2 (in our case).

OpenStudy (mathstudent55):

Can you write an expression of the average of 7 and 5?

OpenStudy (amistre64):

since you divide by 2 in each case, its only the adding thats key

OpenStudy (anonymous):

7+5=12/2=6 e.e

OpenStudy (amistre64):

a/2 = b/2 implies a=b

OpenStudy (mathstudent55):

The expression for the average of 7 and 5 is: \(\dfrac{7 + 5}{2} \) Right?

OpenStudy (anonymous):

yes o.o

OpenStudy (mathstudent55):

The expression of the average of n and 3 is: \(\dfrac{n + 3}{2} \) Since we are told the averages are equal, you set the two expressions equal: \(\dfrac{n + 3}{2}\ = \dfrac{7 + 5}{2} \) Now you need to solve for n. The first step is to add the 7 and 5 in the right numerator. Then multiply both sides by 2. That gets rid of the denominators. Then subtract 3 from both sides to find n.

OpenStudy (anonymous):

Wait...I thought I understood and I was gonna say the answer is C.9 but I'm lost at the last two steps that you just stated...

OpenStudy (phi):

Here is the *important* idea \[ 2 \cdot \frac{1}{2} = \frac{2}{2}= 1\] or if you start with \[ \frac{x}{2} \] you can write that as \[ \frac{1}{2} \cdot x \] now we use the important idea: if we multiply by 2 we get \[ 2 \cdot \frac{1}{2}\cdot x = 1\cdot x = x\]

OpenStudy (phi):

For your problem, \[ \dfrac{n + 3}{2}\ = \dfrac{7 + 5}{2} \] you can "get rid" of the 2 in the bottom by multiplying by 2 up top (this is the important idea) but with an equation, to keep things equal, you have to do the same thing to both sides. So multiply both sides by 2: \[ 2\cdot \dfrac{n + 3}{2}\ = \dfrac{7 + 5}{2} \cdot 2\] notice we have \[ \frac{2}{2} \cdot (n+3) = (7+5) \frac{2}{2} \] (the 2/2 is 1 , which is why we multiplied by 2... it makes 2/2, and that is 1) in other words: n+3 = 7+5 now solve for n the next step is add -3 to both sides. then simplify.